OpenStudy (anonymous):

Simplify the following and write your answer in simplest form. (5/3)^2. A. 25/9 B. 2 7/9 C. 1 2/3 D. 9/25 @texaschic101

OpenStudy (anonymous):

I know you have to do 5/3 x 5/3 and get 25/ 9 but what is the simplest form of it?

OpenStudy (anonymous):

answer is B i just did it

OpenStudy (anonymous):

how is B the answer? what did you do?

OpenStudy (anonymous):

well I did 5 divided by 3^2

OpenStudy (anonymous):

so 5 divided by 9?

OpenStudy (anonymous):

@Somy can you help me explain this lol

OpenStudy (somy):

(5/3)^2 meaning 5^2/3^2 and = 25/9

OpenStudy (anonymous):

this gets me a little tricky the eay i tried to do it was in a different way

OpenStudy (somy):

which is A

OpenStudy (anonymous):

thanks somy

OpenStudy (anonymous):

@iPwnBunnies

OpenStudy (anonymous):

is A the right answer?

OpenStudy (anonymous):

It should be

OpenStudy (ipwnbunnies):

You can break it down to a mixed fraction, but it's arguable to say if that's the simplest form.

OpenStudy (somy):

yes thats right, but this is the only option here @iPwnBunnies

OpenStudy (ipwnbunnies):

What. o.o 2 and 7/9 is the mixed fraction for 25/9. .-. I'm confused now.

OpenStudy (anonymous):

So am I.... O.O

OpenStudy (somy):

wait its 2 7/9???

OpenStudy (somy):

coz if its like this then this one is the right answer i thought it 27/9

OpenStudy (somy):

sadly OS doesn't have the option of writing \[2 \frac{ 7 }{ 9 }\] in the question B is your answer my bad i saw it wrong

OpenStudy (somy):

i mean in questions*

OpenStudy (anonymous):

thanks somy :D

OpenStudy (somy):

you are welcome @Awesome=ME